# Text2AC-Zero: Consistent Synthesis of Animated Characters using 2D Diffusion

### v1

@@ 4.4 Ablation Study
$H_{MSE}\downarrow$
Baseline (Text2Video-Zero [13])
84.87
with SLA
78.90 (-7.0 %)
with PWG
82.73 (-2.5 %)
with LSA + PWG
76.41 (-10.0 %)
Table 2: An ablation study for different components of our proposed approach. SLA: Spatial Latent Alignment in Section 3.4, PWG: Pixel-Wise Guidance in Section 3.5.
@@ 4.5 Limitations and Failure Cases
Since we employ ControlNet with depth conditioning for generating the video frames, our approach is also bounded by its limitations. For example, the top row of Figure 7 shows an example where ControlNet fails to produce a realistic left arm and leg when they intersect in the depth map. Another source of failure is mismatches when computing the correspondence mapping in Section 3.3. These mismatches can lead to inconsistent texture, as shown in the second row of Figure 7. Despite that the Pixel-Wise Guidance module in Section 3.5 attempts to fix these discrepancies, sometimes it does not fully fix them due to insufficient scaling factor in Equation 8. It is also worth mentioning that Pixel-Wise Guidance imposes high GPU memory usage due to computing gradients with respect to the latent codes. However, employing the Spatial Latent Alignment solely still achieves remarkable improvement over the baseline as shown in Table 2 with no GPU memory overhead.
@@
Refer to caption Refer to caption
Figure 7: Examples of failure cases.
@@ 5 Conclusion and Future Work
We introduced a new paradigm for generating consistent videos of animated characters in a zero-shot manner. We employed text-based motion diffusion models to provide continuous motion guidance that we utilized to generate video frames through a pre-trained T2I diffusion model. This allowed generating videos of diverse characters and motions that existing T2V methods struggled to produce. We also demonstrated that our approach produces temporally consistent videos achieved through the proposed Spatial Latent Alignment and Pixel-Wide Guidance modules. We believe that these modules can benefit other latent diffusion-based approaches. For future work, the cross-frame dense correspondences can be improved for better latent alignment. Furthermore, video dynamics can be incorporated into the background for enhanced realism.

### v2

@@ 4.4 Ablation Study
$H_{MSE}\downarrow$ Runtime (s)
Baseline 84.87 28
with SLA 78.90 (-7.0 %) 30
with PWG 82.73 (-2.5 %) 49
with LSA + PWG 76.41 (-10.0 %) 50
Table 2: An ablation study for different components of our proposed approach. SLA: Spatial Latent Alignment in Section 3.4, PWG: Pixel-Wise Guidance in Section 3.5. Runtime is reported for generating a 8-frames video.
@@ 4.5 Limitations and Failure Cases
Since we employ ControlNet with depth conditioning for generating the video frames, our approach is also bounded by its limitations. For example, the top row of Figure 7 shows an example where ControlNet fails to produce a realistic left arm and leg when they intersect in the depth map. Another source of failure is mismatches when computing the correspondence mapping in Section 3.3, which can lead to some artifacts. It is also worth mentioning that Pixel-Wise Guidance imposes high GPU memory usage due to computing gradients with respect to the latent codes. However, employing the Spatial Latent Alignment solely still achieves remarkable improvement over the baseline as shown in Table 2 with no GPU memory overhead.
@@
Figure 7: Examples of a failure case
@@ 5 Conclusion and Future Work
We introduced a new paradigm for generating consistent videos of animated characters in a zero-shot manner. We employed text-based motion diffusion models to provide continuous motion guidance that we utilized to generate video frames through a pre-trained T2I diffusion model. This allowed generating videos of diverse characters and motions that existing T2V methods struggled to produce. We also demonstrated that our approach produces temporally consistent videos achieved through the proposed Spatial Latent Alignment and Pixel-Wide Guidance modules. These two modules can benefit other approaches that adopt cross-frame attention and latent diffusion models in general. For future work, the cross-frame dense correspondences can be improved for better latent alignment. Furthermore, video dynamics can be incorporated into the background for enhanced realism.
@@ 4 Qualitative Ablation Study
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption